what is 7 1/2 + 6 5/8
Answer by Alan3354(69443)

what is 7 1/2 + 6 5/8
-------
Change to improper fractions
= 15/2 + 53/8
Cross multiply to get the numerator
NUM = 15*8 + 2*53
The denominator is the product of the 2 given DENs
DEN = 2*8
--> 226/16
= 113/8